Please help me!!!! I need to solve this for H:
2A-ha=hb
A)h=a-b/2A
B)h=2A/a-b
C)h=a+b/2A
D)h=2A/a+b
Please show me step by step how to figure this out...thanks

Answer by vidhyak(80)

2A-ha=hb
First bring all the "h" terms to one side
2A = hb+ha
Rewriting the equation for better reading
hb+ha = 2A
h(b+a) = 2A
h = 2A/(b+a)

Answer is D